Draw a sketch of the target area and show your calculations for each problem below. Assume that IQ scores are normally distributed with a mean of 100 and a standard deviation of 15. Hint: Use the Empirical Rule (68-95-99.7.)
3. What percent of IQs are below 70?

Answer by stanbon(75887)

Draw a sketch of the target area and show your calculations for each problem below. Assume that IQ scores are normally distributed with a mean of 100 and a standard deviation of 15. Hint: Use the Empirical Rule (68-95-99.7.)
3. What percent of IQs are below 70?
---------
70 = 100 - 2*15
So 70 is 2 standard deviations below the mean.
-------
Since 99.7% are within 2 std's of the mean,
the remaining %age to the left of 70 is (100-99.7)/2 = 0.3/2 = 0.15%
================================
Cheers,
Stan H.